Orders from "heavy user" customers generate about $136 in contribution margin per year for Golf Guide. Golf Guide budgets $53 per year in retention spending per "heavy user" customer. Annual retention rate for these customers is 67%. Customer acquisition cost for these customers is $65 per customer. Assuming a corporate discount rate of 11%, find the customer lifetime value for one "heavy user" customer at Golf Guide. (Rounding: nearest penny.)

Answer: 61.39 (Please show a step-by-step explanation on how to get to this answer. Please do not submit anything if you do not get 61.39 as your answer).
Answer #1

Customer Lifetime Value is a metric to ascertain the present value of the future cash flows of the customer over the lifetime of its relationship with the company. Think of it like NPV of an asset which is acquired at the cost called Acquisition Cost.

think of acquisition cost like an initial investment to acquire the customer

Cash flow is counted as present value of the profits generated by the customer.

Annual Profits = Annual Revenue - Annual Expenses

CLV = future cash flows - acquisition cost

for 1 year, Cash flow = Profit Margin * Retention rate / (1+Discount Rate - Retention rate)

Profit Margin = $136 - $53 = $83

CLV = 83*0.67/(1+0.11-0.67) - 65

CLV = 126.39 - 65 = $ 61.39